# Mirror and Preconditioned Gradient Descent in Wasserstein Space

### v1

@@ I.2 Convergence lemma
We first provide a Lemma which follows from [81, Theorem 3.1], and which is useful for the proofs of Propositions 3, 5 and 23.
@@ I.3 Some properties of Bregman divergences
We provide in this Section additional results on the Bregman divergences introduced in Section 3. First, we focus on $\phi_{\mu}(\mathrm{T})=\int V\circ\mathrm{T}\ \mathrm{d}\mu$. The following Lemma is akin to [73, Proposition 4] which shows it only for OT maps.
@@
[…]
Next, we focus on $\phi_{\mu}(\mathrm{T})=\frac{1}{2}\iint W\big{(}\mathrm{T}(x)-\mathrm{T}(x^{\prime})\big{)}\ \mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime})$, and we generalize the result from [73, Proposition 4].
@@
$\mathrm{d}_{\phi_{\mu}}(\mathrm{T},\mathrm{S})$ $=\phi_{\mu}(\mathrm{T})-\phi_{\mu}(\mathrm{S})-\langle(\nabla W\star\mathrm{S}_{\#}\mu)\circ\mathrm{S},\mathrm{T}-\mathrm{S}\rangle_{L^{2}(\mu)}$ (205)
@@
$\quad-\int\langle(\nabla W\star\mathrm{S}_{\#}\mu)(\mathrm{S}(x)),\mathrm{T}(x)-\mathrm{S}(x)\rangle\ \mathrm{d}\mu(x).$
@@
$=\frac{1}{2}\iint\mathrm{d}_{W}\big{(}\mathrm{T}(x)-\mathrm{T}(x^{\prime}),\mathrm{S}(x)-\mathrm{S}(x^{\prime})\big{)}\ \mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime}).$
@@
$\mathrm{d}_{\phi_{\mu}}(\mathrm{T},\mathrm{S})$ $=\phi_{\mu}(\mathrm{T})-\phi_{\mu}(\mathrm{S})-\langle\nabla_{\mathrm{W}_{2}}\phi(\mathrm{S}_{\#}\mu)\circ\mathrm{S},\mathrm{T}-\mathrm{S}\rangle_{L^{2}(\mu)}$ (209)
@@
$\quad-\langle\mathrm{T}_{\mathrm{S}_{\#}\mu}^{\rho}\circ\mathrm{S}-\mathrm{S},\mathrm{T}_{\mathrm{S}_{\#}\mu}^{\rho}\circ\mathrm{S}-\mathrm{T}_{\mathrm{T}_{\#}\mu}^{\rho}\circ\mathrm{T}\rangle_{L^{2}(\mu)}$
@@
$\quad+\langle\mathrm{T}_{\mathrm{S}_{\#}\mu}^{\rho}\circ\mathrm{S}-\mathrm{S},\mathrm{T}_{\mathrm{T}_{\#}\mu}^{\rho}\circ\mathrm{T}-\mathrm{T}_{\mathrm{S}_{\#}\mu}^{\rho}\circ\mathrm{S}\rangle_{L^{2}(\mu)}.$

### v2

@@ I.2 Convergence lemma
We first provide a Lemma which follows from [88, Theorem 3.1], and which is useful for the proofs of Propositions 4, 6 and 25.
@@ I.3 Some properties of Bregman divergences
We provide in this Section additional results on the Bregman divergences introduced in Section 3. First, we focus on $\phi_{\mu}(\mathrm{T})=\int V\circ\mathrm{T}\ \mathrm{d}\mu$. The following Lemma is akin to [80, Proposition 4] which shows it only for OT maps.
@@
$\mathrm{d}_{\phi_{\mu}}(\mathrm{T},\mathrm{S})$ $=\phi_{\mu}(\mathrm{T})-\phi_{\mu}(\mathrm{S})-\langle\nabla V\circ\mathrm{S},\mathrm{T}-\mathrm{S}\rangle_{L^{2}(\mu)}$
$=\int V\circ\mathrm{T}-V\circ\mathrm{S}-\langle\nabla V\circ\mathrm{S},\mathrm{T}-\mathrm{S}\rangle\ \mathrm{d}\mu$ (209)
$=\int\mathrm{d}_{V}\big{(}\mathrm{T}(x),\mathrm{S}(x)\big{)}\ \mathrm{d}\mu(x).\qed$
Next, we focus on $\phi_{\mu}(\mathrm{T})=\frac{1}{2}\iint W\big{(}\mathrm{T}(x)-\mathrm{T}(x^{\prime})\big{)}\ \mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime})$, and we generalize the result from [80, Proposition 4].
@@
$\mathrm{d}_{\phi_{\mu}}(\mathrm{T},\mathrm{S})$ $=\phi_{\mu}(\mathrm{T})-\phi_{\mu}(\mathrm{S})-\langle(\nabla W\star\mathrm{S}_{\#}\mu)\circ\mathrm{S},\mathrm{T}-\mathrm{S}\rangle_{L^{2}(\mu)}$
@@
$\quad-\int\langle(\nabla W\star\mathrm{S}_{\#}\mu)(\mathrm{S}(x)),\mathrm{T}(x)-\mathrm{S}(x)\rangle\ \mathrm{d}\mu(x).$ (211)
@@
$=\frac{1}{2}\iint\mathrm{d}_{W}\big{(}\mathrm{T}(x)-\mathrm{T}(x^{\prime}),\mathrm{S}(x)-\mathrm{S}(x^{\prime})\big{)}\ \mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime}).\qed$
@@
$\mathrm{d}_{\phi_{\mu}}(\mathrm{T},\mathrm{S})$ $=\phi_{\mu}(\mathrm{T})-\phi_{\mu}(\mathrm{S})-\langle\nabla_{\mathrm{W}_{2}}\phi(\mathrm{S}_{\#}\mu)\circ\mathrm{S},\mathrm{T}-\mathrm{S}\rangle_{L^{2}(\mu)}$
@@
$\quad-\langle\mathrm{T}_{\mathrm{S}_{\#}\mu}^{\rho}\circ\mathrm{S}-\mathrm{S},\mathrm{T}_{\mathrm{S}_{\#}\mu}^{\rho}\circ\mathrm{S}-\mathrm{T}_{\mathrm{T}_{\#}\mu}^{\rho}\circ\mathrm{T}\rangle_{L^{2}(\mu)}$ (215)
@@
$\quad+\langle\mathrm{T}_{\mathrm{S}_{\#}\mu}^{\rho}\circ\mathrm{S}-\mathrm{S},\mathrm{T}_{\mathrm{T}_{\#}\mu}^{\rho}\circ\mathrm{T}-\mathrm{T}_{\mathrm{S}_{\#}\mu}^{\rho}\circ\mathrm{S}\rangle_{L^{2}(\mu)}.\qed$ (216)